An airplane has a speed of 270 km/h with respect to the air. There is a wind blowing at 65 km/h at 30° north of east with respect to Earth. In which direction should the plane head in order to land at an airport due north of its present location?

What would be the plane's speed with respect to the ground?
Answer #1

The east component must be equal and opposite.

x component of wind = 65 * cos 30 = 56.29 Km/h

Therefore,

direction should be

56.29 = 270 * cos x

cos x = 0.20848

x = 77.96 degree

Now, plane's speed with respect to ground

v = 270 * sin 77.96 + 65 * sin 30

v = 264.07 + 32.5

v = 296.57 km/h